OpenStudy (anonymous):

what is the derivative of log 5 sqrt 5x+2

OpenStudy (anonymous):

Log base 5, or base 10?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Ok hang on.

OpenStudy (anonymous):

I was thinking 1/ln5 * (5/2sqrt5x+2) now I'm stuck

OpenStudy (anonymous):

So maybe it is dy/dx= 5/(ln5)(5x+2)

OpenStudy (anonymous):

Ok let u equal sqrt(5x+2). So we have. \[\log_{5}u \] The derivative is then \[\frac{ 1 }{ u \ln 5 }\] In this case 'u' is \[\sqrt{5x+2}\] So we have \[\frac{ 1 }{ \sqrt{5x+2}\ln5 }\]

OpenStudy (anonymous):

Now by the definition of the chain rule we have to multiply by the derivative of \[\sqrt{5x+2}\] and that is \[\frac{ 5 }{ 2\sqrt(5x+2) }\]

OpenStudy (anonymous):

I like where this is going...I think I am on track

OpenStudy (anonymous):

So now it's \[\frac{ 1 }{ \ln5\sqrt(5x+2) } * \frac{ 5 }{ 2\sqrt(5x+2) }\] and that is \[\frac{ 5 }{ \ln5(10x+4) }\] You might have to check that though.

OpenStudy (anonymous):

Sorry, online class so I am teaching myself from the book basically. Why did you multiply?

OpenStudy (anonymous):

O gosh, wow...wow that was stupid. I see

OpenStudy (anonymous):

So the chain rule is The derivative of f(g(x)) is f'(g(x))*g'(x) In this case f(x) is log 5 x and g(x) is sqrt(5x+2) The annoying part is that you have to apply the chain rule again to sqrt(5x+2)

OpenStudy (anonymous):

Alright, sweet

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

np!